OpenStudy (anonymous):

Find the exact values of the sine and cosine of 315˚ and -315˚. Then find the decimal equivalents. Round your answers to the nearest hundredth.

OpenStudy (anonymous):

you got a nice unit circle cheat sheet?

OpenStudy (anonymous):

now you do find \(315^\circ\) on the last page of the attached cheat sheet then look at the corresponding coordinates on the unit circle the first coordinate is cosine, the second is sine

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

\[y=\sin ^{-1}y\] \[x=\sin(y)\]

OpenStudy (anonymous):

you lost me there aren't you trying to find \(\sin(315^\circ)\)?

OpenStudy (anonymous):

i guess. i honestly have no idea what its asking me

OpenStudy (anonymous):

that is what is says right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

"Find the exact values of the sine and cosine of 315˚"

OpenStudy (anonymous):

i don't get the page you posted

OpenStudy (anonymous):

are you on the last page with the circle?

OpenStudy (anonymous):

thats the 2nd to last page

OpenStudy (anonymous):

actually it is page 3, second to last right

OpenStudy (anonymous):

you see 315 labelled on the circle?

OpenStudy (anonymous):

yes now you see that ordered pair there right, looks like \[(\frac{\sqrt2}{2},-\frac{\sqrt2}{2})\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the first coordinate is \(\cos(315^\circ)\) and the second is \(\sin(315^\circ)\)

OpenStudy (anonymous):

so \[\cos(315^\circ)=\frac{\sqrt2}{2}\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

easy enuf right?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

let me know when you need help with \(-315^\circ\)

OpenStudy (triciaal):

to the nearest hundredth means to 2 decimal places